given qr=7ft amd rs is three times the length of qr find the lenghth of qs
1
Expert's answer
2013-03-15T09:42:45-0400
given qr=7ft amd rs is three times the length of qr find thelenghth of qs.rs is three timesthe length of qr:
rs = 3*qr
qr=7ft => rs=3*7ft=21ft
r is between q and s,therefore:
qr+rs=qs
qr=7ft and rs=21ft
qs=7ft+21ft=28ft
Answer: qs=28ft